40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Sep 7 23:23:00 2018, awade, anchal, DailyProgress, BEAT, Back to beat 
    Reply  Wed Sep 12 16:27:19 2018, anchal, DailyProgress, BEAT, Back to beat NorthandSouth_14750kHz_RFPD_Measurements.pdf
       Reply  Fri Sep 14 00:57:14 2018, rana, DailyProgress, BEAT, Back to beat 
       Reply  Fri Sep 21 18:10:53 2018, rana, DailyProgress, BEAT, Back to beat 
Message ID: 2230     Entry time: Fri Sep 7 23:23:00 2018     Reply to this: 2232
Author: awade, anchal 
Type: DailyProgress 
Category: BEAT 
Subject: Back to beat 

[awade, anchal​]

Today we go the beat to converge on the 26 MHz resonance of the resonant transmission beat detector.  Anchal and I took a TF of the PLL and found the UGF to be 40 kHz, sufficient for a BN spectrum.  Beat not power was about -1 dBm.  

Sorry no beat spectrum, we lost lock with the south path autolocker turned off and the cavity heat PID didn't realize and knocked the BN way off.  Going to take another 12 hours for it to settle down again.

It looks like the spectrum was about 0.1 Hz at its lowest point (>3 kHz).  We saw that there was a very prominent narrow peak at about 2.8 kHz.  We had a look at the mixer monitor signals and saw that the narrow peak was coming from the south path FSS.  After doing a cross spectrum between the beat and the north and south FSS it became apparent that a significant portion of our noise is coming from the south FSS somehow.  We need to track down exactly what would be coupling in a 2.8 kHz narrow peak and also the broader noise of the south FSS.  

I suspect​ it might be a photodetector​ issue.  We had excess noise is one path relative to the other before and swapped detectors and field boxes.  We never quite diagnosed it but it could well be that the 14.75 MHz tuned detector that is now in the south path is responsible.  The first thing to check is the TF by injecting into the test port.  Also the dark noise should be checked around 14.75 MHz.  The optical TF can also be taken with the Jenny rig at the 40m.  Now might be a good time to switch modulation frequencies to 36 MHz and 37 MHz.  These are fine for HOM spacing and we have the Wenzel crystals ready to go.  We also have two detectors tuned to 35.5 MHz sitting on the shelf and BB EOM drivers that just need stuffing onto boards.  This might take Anchal a week to do, he seems to be good with electronics.

Another unresolved problem is the residual AM from the 14.75 MHz phase modulators.  I was never really able to reduce this down and keep it down.  Thermal or alignment drift seemed to make it really hard to minimize.  It could be bad alignment though the crystal or thermal drift.  They have insulating hats on but they still are less than optimal.  An ISS will do a bit to suppress noise opened up by this effect but we would like to solve it properly.

ELOG V3.1.3-